# Examples of Ricci limit spaces with infinite holes

### v1

@@ Abstract.
In this paper, we construct, for each given $\epsilon>0$, a sequence of $(n+2)$-dimensional manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ with $\mathrm{Ric}_{g_{i}}>\lambda$, such that $d_{GH}(X,X_{\epsilon})\leq\epsilon$, and $X_{\epsilon}$ is homeomorphic to the space obtained by removing an infinite number of balls from $X$. Hence $X_{\epsilon}$ has dense boundary with an infinite number of connected components. Moreover, $X_{\epsilon}$ has no open subset which is topologically a manifold. This generalizes Hupp-Naber-Wang’s result [11] from $4$-dimensional case to the general case of dimension $n\geq 3$. Notably, the restriction $n\geq 3$ is essential, as Ricci limit spaces in $1$ and $2$ dimensions always have a dense open subset which is topologically a manifold [10, 7, 13].
Our construction differs from that of Hupp-Naber-Wang. In their approach, Hupp-Naber-Wang considered doing an infinite number of blow-ups on the local complex surface structure of $X$, thus relying on the $4$-dimensional condition. However, our method involves removing an infinite number of balls from $X$, allowing us to construct in the general case of dimensions greater than or equal to $3$.
As a corollary, we provide a solution to an open problem posed by Naber [15, Open Problem 3.4] in the $3$-dimensional case. Note that the $3$-dimensional case is the only remaining unresolved case in Naber’s problem. Consequently, Naber’s question is resolved in all dimensions.
@@ 1. Introduction
In the non-collapsed case, which is to say $\mathrm{Vol}(B_{1}(p_{i}))\geq v>0$, the works of Colding [5, Theorem 0.1] and Cheeger-Colding [2, Theorem 5.9] show that the measure $\nu=\mathcal{H}^{n}$ is just the $n$-dimensional Hausdorff measure on $(X^{n},d)$, and the set of $n$-regular points $R_{n}(X)\subset X$ is a subset of full-$\nu$ measure, where a point $x\in X$ is called a $k$-regular point if and only if every tangent cone of $X$ at $x$ is isometric to $\mathbb{R}^{k}$. Then a Reifenberg type result of Cheeger-Colding [2, Theorem 5.14] shows that the limit space has a manifold structure away from a codimension $2$ set, which is the complement of an open neighborhood of $R_{n}(X)$. More recently, the work of Cheeger-Jiang-Naber [4, Theorem 1.14] was able to prove that in this case the codimension $2$ Hausdorff measure of the subset we removed can be finite.
@@
Problem 1.1 ([15, Open Problem 3.4]).
Let $(M_{i},g_{i},\nu_{i},p_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X^{n},d,\nu,p)$ with $\mathrm{Ric}_{g_{i}}\geq-\lambda$ and $\mathrm{Vol}(B_{1}(p_{i}))\to 0$, then is there an open subset of full-$\nu$ measure $R(X)\subset X$ which is homeomorphic to a $n$-manifold?
It’s worth noting that for $n=1,2$, the answer of Problem 1.1 is affirmative [10, 7, 13].
@@
However, Problem 1.1 is still open when $n=3$. On one hand, the blow-up surgery used by Hupp-Naber-Wang requires local complex structures and cannot be applied to $3$-manifolds. On the other hand, Simon-Topping’s result [18] about the manifold structure of $3$-dimensional non-collapsing Ricci limit spaces indicates that there may be some additional topological regularity in the $3$-dimensional case compared to higher dimensions.
In this paper, we construct a sequence of $(n+2)$-dimensional manifolds for any given $n$-dimensional manifolds $(n\geq 3)$ with Ricci lower bound such that the limit space has no open subset which is a topological manifold.
The following theorem is our main theorem in this paper.
@@
    For any open subset $U$ of $X_{\epsilon}$, the homology group $H_{n}(U;\mathbb{Z})$ is infinitely generated. Consequently, every open subset of $X_{\epsilon}$ is not a topological manifold.
@@
Note here that the dimension condition $n\geq 3$ in Theorem 1.1 is interesting. We would like to point out that this condition is sharp. Thus we can refine Hupp-Naber-Wang’s answer to Problem 1.1. Specifically, the following facts are now known:
  • When $n=1$, the Ricci limit space of rectifiable dimension $n$ is always a topological manifold with boundary [10, 7].
  • When $n=2$, the Ricci limit space of rectifiable dimension $n$ is always an Alexandrov space [13], and hence there exists a dense open subset of such space homeomorphic to a topological manifolds.
@@
Now we can define the boundary of $X$ by
(3) $\partial X=\mathcal{S}^{n-1}\backslash\mathcal{S}^{n-2}.$
Here we follow the definition used by Bruè-Naber-Semola [1, Definition 1.1]. It’s important to note that there are other definitions of boundary. For instance, De Philippis-Gigli [8] and Kapovitch-Mondino [12] have provided different definitions of the boundary. The motivation for studying the geometry of boundaries comes from the research about $\mathrm{RCD}$ spaces.
Note that Cheeger-Colding [2, Theorem 6.1] proved that the boundary of non-collapsed Gromov-Hausdorff limits of boundary free manifolds is always empty. As a corollary of Theorem 1.1, our construction also gives examples where boundaries are both dense and contain an infinite number of connected components.
@@
Let $n\geq 3$. Then there exists a sequence of $(n+2)$-dimensional boundary free complete manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X^{n}_{\infty},d_{\infty})$ with $\mathrm{Ric}_{g_{i}}\geq-1$, such that $\partial X_{\infty}$ is both dense and contains an infinite number of connected components.
@@
Remark.
Note that Kapovitch-Mondino used tangent cones to define the boundary, while the boundary defined by De Philippis-Gigli is the closure of the boundary defined by Bruè-Naber-Semola. Thus, one can see that Corollary 1.1 holds for the boundary defined by Kapovitch-Mondino. However, the boundary defined by De Philippis-Gigli on $X_{\infty}$ is $X_{\infty}$.
@@
Acknowledgement. The author wants to express his deep gratitude to Professor Gang Tian for constant encouragement. He would also be grateful to Professor Aaron Naber, Professor Wenshuai Jiang and Professor Jianchun Chu for valuable conversations. Additionally, he expresses gratitude to Professor Xin Fu and Professor Jiyuan Han for their warm hospitality during his visit to Westlake University.
@@ 2. Outline of the construction
The above graph is to provide some intuition. Although we can construct it topologically, the above construction is impossible in the Riemannian geometric sense. We point out again that due to the deep result of Lytchak-Stadler [13], the $2$-dimensional Ricci limit space must be an Alexandrov space. Now we give an outline of our construction.
@@ 3.1. Double warped products
See also [17, Proposition 1.4.7] for more details.
@@
One can conclude it by a straightforward calculation. See also [17, Subsection 4.2.4]. ∎
@@
This is a consequence of Menguy’s $C^{1}$ gluing lemma [14, Lemma 1.170]. See also [11, Lemma 3.1].
@@ A Riemannian metric $g_{X,i,k}$ and a positive function $f_{i,k}$ on $X\backslash\mathcal{W}_{i,k}$,
    For any open subset $U$ of $X_{\epsilon}$, the homology group $H_{n}(U;\mathbb{Z})$ is infinitely generated. Consequently, every open set of $X_{\epsilon}$ is not a topological manifold.
@@
Now we consider the property (d). Let $U$ be an open subset of $X_{\epsilon}$. Assume that the homology group $H_{n}(U)$ is finitely generated. Without loss of generality, we can assume that $\mathrm{rank}(H_{n}(U))=N_{U}\in\mathbb{N}$.
@@
Now we recall that the inclusion map $I_{i,i-1}:X\backslash\bar{\mathcal{W}}_{i}\to X\backslash\bar{\mathcal{W}}_{i-1}$ is an $(1+2^{-i}\epsilon)$-Lipschitz map, and induces an $(1+2^{-i}\epsilon)$-bi-Lipschitz map on the connected components $\partial\mathcal{U}_{j}\to\partial\mathcal{U}_{j}$ of the boundary of $\partial\mathcal{W}_{i}$. Then we can see that the inclusion maps $I_{i_{N}+1,i_{N}}\circ\cdots\circ I_{i-1,i-2}\circ I_{i,i-1}$ converging to a Lipschitz map $X_{\epsilon}\to X\backslash\mathcal{W}_{i_{N}}$, and induces a bi-Lipschitz map between $\partial\mathcal{W}_{i_{N}}$ and a compact subset $\mathcal{Y}_{i_{N}}^{\infty}$ of $X_{\epsilon}$. Hence for $j=1,\cdots,N$, the image of $H_{n}(U)\to H_{n}(X\backslash\mathcal{W}_{i_{N}})$ induced by $U\subset X_{\epsilon}\to X\backslash\mathcal{W}_{i_{N}}$ contains the image of $\bigoplus_{j=1}^{N}H_{n}(\partial\mathcal{U}_{j})=H_{n}(\mathcal{W}_{i_{N}})\to H_{n}(X\backslash\mathcal{W}_{i_{N}})$ induced by the inclusion map $\bigsqcup_{j=1}^{N}\partial\mathcal{U}_{j}=\mathcal{W}_{i_{N}}\to X\backslash\mathcal{W}_{i_{N}}$. On the other hand, we have
$\mathrm{rank}\left(\mathrm{image}(H_{n}(U)\to H_{n}(X\backslash\mathcal{W}_{i_{N}}))\right)\leq\mathrm{rank}(H_{n}(U))\leq N_{U}.$
@@
$\mathrm{rank}\left(\mathrm{image}\left(\bigoplus_{j=1}^{N}H_{n}(\partial\mathcal{U}_{j})\to H_{n}(X\backslash\mathcal{W}_{i_{N}})\right)\right)\leq\mathrm{rank}\left(\mathrm{image}(H_{n}(U)\to H_{n}(X\backslash\mathcal{W}_{i_{N}}))\right)\leq N_{U}.$
@@
$\mathrm{rank}\left(\mathrm{image}\left(\bigoplus_{j=1}^{N}H_{n}(\partial\mathcal{U}_{j})\to H_{n}((X/\thicksim)\backslash\mathcal{W}_{i_{N}})\right)\right)\leq N_{U}.$
@@
$0=H_{n+1}(X/\thicksim)\to\bigoplus_{j=1}^{N}H_{n}(\partial\mathcal{U}_{j})\to H_{n}((X/\thicksim)\backslash\mathcal{W}_{i_{N}})\bigoplus H_{n}\left(\cup_{j=1}^{N}\bar{\mathcal{U}}_{j}\right)\to H_{n}(X/\thicksim).$
@@
$N=\mathrm{rank}\left(\bigoplus_{j=1}^{N}H_{n}(\partial\mathcal{U}_{j})\right)\leq\mathrm{rank}\left(\mathrm{image}\left(\bigoplus_{j=1}^{N}H_{n}(\partial\mathcal{U}_{j})\to H_{n}((X/\thicksim)\backslash\mathcal{W}_{i_{N}})\right)\right)\leq N_{U}.$
@@
Proof of Corollary 1.1: Let $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ be the sequence of $(n+2)$-dimensional manifolds we constructed in Proposition 5.1. Then Proposition 5.1 implies that for any $x\in\mathcal{Y}_{j}^{\infty}$, all tangent cone at $x$ are isometric to $\mathbb{R}^{n-1}\times\mathbb{R}_{\geq 0}$, where $\mathcal{Y}_{j}^{\infty}$ is the limit of $\mathcal{W}_{j}$ in $X_{\epsilon}$. Hence $\cup_{j=1}^{\infty}\mathcal{Y}_{j}^{\infty}\subset\partial X_{\epsilon}$ is dense in $X_{\epsilon}$. Applying Proposition 5.1 again, one can conclude that $\cup_{i=1}^{\infty}\partial B^{h}_{2^{-i}}\subset R_{n}(p_{0})$. Then $\cup_{i=1}^{\infty}(\partial B^{h}_{2^{-i}}\cap\partial X_{\epsilon})=\emptyset$, and hence $\partial X_{\epsilon}$ has an infinite number of connected components. ∎

### v2

@@ Abstract.
In this paper, we construct, for each given $\epsilon>0$, a sequence of $(n+2)$-dimensional manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ with $\mathrm{Ric}_{g_{i}}>\lambda$, such that $d_{GH}(X,X_{\epsilon})\leq\epsilon$, and $X_{\epsilon}$ is homeomorphic to the space obtained by removing an infinite number of balls from $X$. Hence $X_{\epsilon}$ has dense boundary with an infinite number of connected components. Moreover, $X_{\epsilon}$ has no open subset which is topologically a manifold. This generalizes Hupp-Naber-Wang’s result [11] from $4$-dimensional case to the case of dimension $n\geq 3$.
Our construction differs from that of Hupp-Naber-Wang. In their approach, Hupp-Naber-Wang considered doing an infinite number of blow-ups on the local complex surface structure of $X$, thus relying on the $4$-dimensional condition. However, our approach avoids the use of local complex surface structure of $X$, allowing us to construct in the general case of dimensions greater than or equal to $3$. As a corollary, we provide a solution to an open problem posed by Naber [16, Open Problem 3.4] in the $3$-dimensional case.
@@ 1. Introduction
In the non-collapsed case, which is to say $\mathrm{Vol}(B_{1}(p_{i}))\geq v>0$, the works of Colding [5, Theorem 0.1] and Cheeger-Colding [2, Theorem 5.9] show that the measure $\nu=c\mathcal{H}^{n}$ for some $c>0$, and the set of $n$-regular points $R_{n}(X)\subset X$ is a subset of full-$\nu$ measure, where $\mathcal{H}^{n}$ is the $n$-dimensional Hausdorff measure on $(X^{n},d)$, and a point $x\in X$ is called a $k$-regular point if and only if every tangent cone of $X$ at $x$ is isometric to $\mathbb{R}^{k}$. Then a Reifenberg type result of Cheeger-Colding [2, Theorem 5.14] shows that the limit space has a manifold structure away from a codimension $2$ set, which is the complement of an open neighborhood of $R_{n}(X)$. More recently, the work of Cheeger-Jiang-Naber [4, Theorem 1.14] was able to prove that in this case the codimension $2$ Hausdorff measure of the subset we removed can be finite.
@@
Problem 1.1 ([16, Open Problem 3.4]).
Let $(M_{i},g_{i},\nu_{i},p_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X^{n},d,\nu,p)$ with $\mathrm{Ric}_{g_{i}}\geq-\lambda$ and $\mathrm{Vol}(B_{1}(p_{i}))\to 0$, then is there an open subset of full-$\nu$ measure $R(X)\subset X$ which is homeomorphic to an $n$-manifold?
It’s worth noting that for $n=1$, the answer of Problem 1.1 is affirmative [10, 7].
@@
However, Problem 1.1 is still open when $n=2$ or $3$.
On one hand, the blow-up surgery used by Hupp-Naber-Wang requires local complex surface structures and cannot be applied to $2$ or $3$-manifolds. On the other hand, Simon-Topping’s result [19], about the manifold structure of $3$-dimensional non-collapsing Ricci limit spaces indicates that there may be some additional topological regularity in the $3$-dimensional case compared to higher dimensions. Here we pay special attention to the $3$-dimensional case, because in many cases people tend to focus on the geometry of $6$-manifolds collapse to a $3$-space, such as the famous metric Strominger-Yau-Zaslow conjecture [13].
In this paper, we construct a sequence of $(n+2)$-dimensional manifolds for any given $n$-dimensional manifolds $(n\geq 3)$ with Ricci lower bound such that the limit space has no open subset which is a topological manifold. The following theorem is our main theorem in this paper.
@@
    For any open subset $U$ of $X_{\epsilon}$, the homology group $H_{n-1}(U;\mathbb{Z})$ is infinitely generated. Consequently, every open subset of $X_{\epsilon}$ is not a topological manifold.
@@
Now we can refine Hupp-Naber-Wang’s answer to Problem 1.1. Specifically, the following facts are now known:
  • When $n=1$, the Ricci limit space of rectifiable dimension $n$ is always a topological manifold with boundary. For more details, see Honda [10] or Chen [10].
@@
Now we can define the reduced boundary of $X$ by
(3) $\partial^{*}X=\mathcal{S}^{n-1}\backslash\mathcal{S}^{n-2}.$
There are two definitions of boundary. The definition of boundary given by De Philippis-Gigli [8], $\partial_{\rm DPG}X$, is just the closure of the reduced boundary. The other definition of a boundary, provided by Kapovitch-Mondino [12], $\partial_{\rm KM}X$, is defined by induction on tangent cones. See also [1] for more details about boundaries.
The motivation for studying the geometry of boundaries comes from the research about $\mathrm{RCD}$ spaces. Note that Cheeger-Colding [2, Theorem 6.1] proved that the boundary of non-collapsed Gromov-Hausdorff limits of boundary free manifolds is always empty. As a corollary of Theorem 1.1, our construction also gives examples where boundaries are both dense and contain an infinite number of connected components.
@@
Let $n\geq 3$. Then there exists a sequence of $(n+2)$-dimensional boundary free complete manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X^{n}_{\infty},d_{\infty})$ with $\mathrm{Ric}_{g_{i}}\geq-1$, such that $\partial_{KM}X_{\infty}$ is both dense and contains an infinite number of connected components. Moreover, $\partial_{\rm DPG}X_{\infty}=X_{\infty}$.
@@
Acknowledgement. The author wants to express his deep gratitude to Professor Gang Tian for constant encouragement. The author would also be grateful to Professor Aaron Naber, Professor Wenshuai Jiang and Professor Jianchun Chu for valuable conversations. Additionally, the author thanks Professor Xiaochun Rong and Professor Guofang Wei for their interests in this note. Furthermore, the author expresses gratitude to Professor Xin Fu and Professor Jiyuan Han for their warm hospitality during his visit to Westlake University. In particular, the author would like to thank Daniele Semola for pointing out some errors in Introduction of the first edition.
@@ 2. Outline of the construction
The above graph is to provide some intuition. Now we give an outline of our construction.
@@ 3.1. Double warped products
See also [18, Proposition 1.4.7] for more details.
@@
One can conclude it by a straightforward calculation. See also [18, Subsection 4.2.4]. ∎
@@
This is a consequence of Menguy’s $C^{1}$ gluing lemma [15, Lemma 1.170]. See also [11, Lemma 3.1].
@@ A Riemannian metric $g_{X,i,k}$ and a positive function $f_{i,k}$ on $X\backslash\mathcal{W}_{i,k}$,
    For any open subset $U$ of $X_{\epsilon}$, the homology group $H_{n-1}(U;\mathbb{Z})$ is infinitely generated. Consequently, every open set of $X_{\epsilon}$ is not a topological manifold.
@@
Now we consider the property (d). Let $U$ be an open subset of $X_{\epsilon}$. Assume that the homology group $H_{n}(U)$ is finitely generated. Without loss of generality, we can assume that $\mathrm{rank}(H_{n-1}(U))=N_{U}\in\mathbb{N}$.
@@
Now we recall that the inclusion map $I_{i,i-1}:X\backslash\bar{\mathcal{W}}_{i}\to X\backslash\bar{\mathcal{W}}_{i-1}$ is an $(1+2^{-i}\epsilon)$-Lipschitz map, and induces an $(1+2^{-i}\epsilon)$-bi-Lipschitz map on the connected components $\partial\mathcal{U}_{j}\to\partial\mathcal{U}_{j}$ of the boundary of $\partial\mathcal{W}_{i}$. Then we can see that the inclusion maps $I_{i_{N}+1,i_{N}}\circ\cdots\circ I_{i-1,i-2}\circ I_{i,i-1}$ converging to a Lipschitz map $X_{\epsilon}\to X\backslash\mathcal{W}_{i_{N}}$, and induces a bi-Lipschitz map between $\partial\mathcal{W}_{i_{N}}$ and a compact subset $\mathcal{Y}_{i_{N}}^{\infty}$ of $X_{\epsilon}$. Hence for $j=1,\cdots,N$, the image of $H_{n-1}(U)\to H_{n-1}(X\backslash\mathcal{W}_{i_{N}})$ induced by $U\subset X_{\epsilon}\to X\backslash\mathcal{W}_{i_{N}}$ contains the image of $\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})=H_{n-1}(\mathcal{W}_{i_{N}})\to H_{n-1}(X\backslash\mathcal{W}_{i_{N}})$ induced by the inclusion map $\bigsqcup_{j=1}^{N}\partial\mathcal{U}_{j}=\mathcal{W}_{i_{N}}\to X\backslash\mathcal{W}_{i_{N}}$. On the other hand, we have
$\mathrm{rank}\left(\mathrm{image}(H_{n-1}(U)\to H_{n-1}(X\backslash\mathcal{W}_{i_{N}}))\right)\leq\mathrm{rank}(H_{n-1}(U))\leq N_{U}.$
@@
$\mathrm{rank}\left(\mathrm{image}\left(\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})\to H_{n-1}(X\backslash\mathcal{W}_{i_{N}})\right)\right)\leq\mathrm{rank}\left(\mathrm{image}(H_{n-1}(U)\to H_{n-1}(X\backslash\mathcal{W}_{i_{N}}))\right)\leq N_{U}.$
@@
$\mathrm{rank}\left(\mathrm{image}\left(\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})\to H_{n-1}((X/\thicksim)\backslash\mathcal{W}_{i_{N}})\right)\right)\leq N_{U}.$
@@
$H_{n}(X/\thicksim)\to\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})\to H_{n-1}((X/\thicksim)\backslash\mathcal{W}_{i_{N}})\bigoplus H_{n-1}\left(\cup_{j=1}^{N}\bar{\mathcal{U}}_{j}\right)\to H_{n-1}(X/\thicksim).$
@@
$N=\mathrm{rank}\left(\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})\right)$$\leq$ $\mathrm{rank}\left(\mathrm{image}\left(\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})\to H_{n-1}((X/\thicksim)\backslash\mathcal{W}_{i_{N}})\right)\right)+\mathrm{rank}H_{n}(X/\thicksim)$
$\leq$ $N_{U}+\mathrm{rank}H_{n}(X/\thicksim)$
@@
Proof of Corollary 1.1: Let $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ be the sequence of $(n+2)$-dimensional manifolds we constructed in Proposition 5.1. Then Proposition 5.1 implies that for any $x\in\mathcal{Y}_{j}^{\infty}$, all tangent cone at $x$ are isometric to $\mathbb{R}^{n-1}\times\mathbb{R}_{\geq 0}$, where $\mathcal{Y}_{j}^{\infty}$ is the limit of $\mathcal{W}_{j}$ in $X_{\epsilon}$. Hence $\cup_{j=1}^{\infty}\mathcal{Y}_{j}^{\infty}\subset\partial X_{\epsilon}$ is dense in $X_{\epsilon}$. Applying Proposition 5.1 again, one can conclude that $\cup_{i=1}^{\infty}\partial B^{h}_{2^{-i}}\subset R_{n}(p_{0})$. Then $\cup_{i=1}^{\infty}(\partial B^{h}_{2^{-i}}\cap\partial X_{\epsilon})=\emptyset$, and hence $\partial X_{\epsilon}$ has an infinite number of connected components. Thus $\partial_{\rm DPG}X_{\epsilon}=X_{\epsilon}$. Similarly, $\partial_{\rm KM}X_{\epsilon}$ has an infinite number of connected components. ∎